# O45 O1em m #3

### v1

@@ 5.4 Ablation Study
5.5 Visualization
@@ 6 Conclusion
We introduce AG-VPReID, a comprehensive benchmark dataset for video-based aerial-ground person ReID, addressing the critical need for a large and challenging aerial-ground dataset. We propose AG-VPReID-Net, a purpose-built three-stream person ReID framework that combines temporal-spatial processing, physics-informed normalized appearance representation, and multi-scale attention mechanisms. Our approach demonstrates state-of-the-art performance on both the AG-VPReID dataset and existing video-based ReID benchmarks. The lower performance of all approaches on the AG-VPReID dataset underscores its challenging nature and provides a robust benchmark for future research.

### v2

@@ 5.5 Visualization
Table 7: Impact of clothing changes (CC) vs. camera angles.
@@ 6 Conclusion
We introduce AG-VPReID, a comprehensive dataset for video-based aerial-ground person ReID, addressing the critical need for a large and challenging aerial-ground dataset. We also propose AG-VPReID-Net, a purpose-built three-stream person ReID framework that combines temporal-spatial processing, physics-informed normalized appearance representation, and multi-scale attention mechanisms. This approach achieves state-of-the-art performance on both the AG-VPReID dataset and existing video-based ReID benchmarks. Notably, the relatively lower performance across all approaches on AG-VPReID highlights its demanding nature and establishes it as a robust benchmark for advancing future research in the field.